OpenStudy (anonymous):

Show all work in dividing 4t^2-1 over 15t+10 divided by 4t^2-4t+1 over 33t^2+5t+2

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

imma senior y?

OpenStudy (anonymous):

this sems like 12? im a ninethgrader but taking ap class let me see if i can do this

OpenStudy (anonymous):

ok thanx

OpenStudy (anonymous):

they are seperated and thenyou have to divide two problems correct? is this similar of what you gbeen looking in your clas?

OpenStudy (anonymous):

\[\frac{4t^2-1}{15t+10}\times \frac{33t^2+5t+2}{4t^2-4t+1}\] is a start

OpenStudy (anonymous):

then you want to factor each term if you can, and cancel any common factors

OpenStudy (anonymous):

yeah that is how i have it

OpenStudy (anonymous):

i no how to start it but idk how to end it .

OpenStudy (anonymous):

@satellite73 can you simplyfy it to see the result please if you

OpenStudy (anonymous):

cece show your simplyfication what you did afterwards

OpenStudy (anonymous):

\[4t^2-1=(2t+1)(2t-1)\] \[33t^2+5t+2\] does not factor, is there a typo there ?

OpenStudy (anonymous):

idk lemmi see.

OpenStudy (anonymous):

\[15t+10=5(3t+2)\] and \[4t^2-4t+1=(2t-1)(2t-1)\]

OpenStudy (anonymous):

i did have a typo its not 33t^2 its 3t^2

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

i am going to guess that it is \[3t^2+5t+2\] in which case you get \[3t^2+5t+2=(t+1)(3t+2)\]

OpenStudy (anonymous):

put it all together and cancel away

OpenStudy (anonymous):

k thanx:)